# PooDLe[Uncaptioned image]: Pooled and dense self-supervised learning from naturalistic videos

### v1

@@ Abstract
Self-supervised learning has driven significant progress in learning from single-subject, iconic images. However, there are still unanswered questions about the use of minimally-curated, naturalistic video data, which contain dense scenes with many independent objects, imbalanced class distributions, and varying object sizes. In this paper, we propose a novel approach that combines an invariance-based SSL objective on pooled representations with a dense SSL objective that enforces equivariance to optical flow warping. Our findings indicate that a unified objective applied at multiple feature scales is essential for learning effective image representations from high-resolution, naturalistic videos. We validate our approach on the BDD100K driving video dataset and the Walking Tours first-person video dataset, demonstrating its ability to capture spatial understanding from a dense objective and semantic understanding via a pooled representation objective.
@@ 1 Introduction
Humans and other animals learn visual understanding from a continuous stream of inputs with little explicit supervision. Recently, self-supervised learning (SSL) [11, 20, 12, 7, 3, 26, 2, 4, 25] has made great strides in learning without human annotations, becoming competitive with supervised learning. However, most methods still revolve around ImageNet [15] which is implicitly supervised through iconic images that contain an unambiguous subject sampled from a balanced class distribution. In contrast, naturalistic vision data such as egocentric videos, contain cluttered scenes, imbalanced classes and objects of varying size, making them ill-suited for existing iconic methods.
@@
Nevertheless, these naturalistic videos are still valuable for their information density and ease of collection, while also mimicking the real-life perspective of humans. Unfortunately, the aforementioned iconic methods, which learn to align pooled representations of an image, may perform poorly as dense scenes often produce views depicting unrelated subjects that are semantically incompatible (Figure 1, red boxes). Recent works have attempted to address this weakness by introducing mechanisms to account for multiple subjects, such as new cropping [39] or attention [42, 35] strategies. Others [49, 45] have also proposed variations using a “dense SSL” objective, which does not pool image representations and instead identifies correspondences between views before computing an SSL loss for each corresponding region.
@@
While dense SSL methods avoid the problem of aligning semantically incompatible views, and can learn fine-grained object boundaries [49], they are susceptible to other issues. Computing the loss at each spatial location presents a spatial imbalance problem, where larger background classes such as the sky dominate the representation while smaller classes such as pedestrians are underrepresented. This effect is undesirable because foreground objects, which are often smaller, should be favored over low-detail, repetitive background classes which often occupy more space. Furthermore, ignoring small foreground objects can be acutely hazardous in applications like self-driving [50] where critical objects, such as traffic lights and pedestrians occupy less than $0.3\%$ of a video frame (Figure 1, green boxes and 1). Preserving these crucial details is particularly necessary in the SSL, where there are no labels to assign importance. This is in contrast to ImageNet [15] training, where models that can easily learn semantics from iconic data with clear, single-subject views and a balanced class distribution. Some dense SSL methods [42, 35, 45] also include losses that optimize a global, pooled representation to learn semantic information from dense scenes, but do not explore how to integrate these two objectives through means of architecture and augmentation strategies.
To address problems of cluttered scenes and spatial imbalance when learning from dense, naturalistic videos, we propose a joint Pooled and Dense Learning (PooDLe) method that optimizes a dense SSL objective over representations without pooling and a secondary pooled objective applied to smaller, content-aligned views. The combination of dense and pooled objectives captures both scene-level understanding and fine-grained boundaries, as well as high-level semantics and small objects. Like FlowE [49], our dense objective encourages the alignment of dense feature maps using optical flow warping. We also use a flow-informed cropping procedure to generate pairs of smaller views, or “subcrops”, with improved subject correspondence between each subcrop pair when training the pooled loss. Depicted in Figure 1, these subcrops often serve as pseudo-iconic views of foreground objects, which functionally increase the prevalence of smaller foreground classes in our dataset. We also introduce a lightweight spatial decoder module (SDM), comprised of top-down decoder layers and UNet [38]-like lateral connections to upsample high-level semantic representations to preserve smaller objects in our dense objective. We find the inclusion of both objectives along with the SDM is essential to capture the semantics of smaller objects and to achieve overall strong downstream task performance.
We conduct experiments on the BDD100K [50] dataset of dashcam driving videos and the recent WalkingTours dataset of first-person walking videos [42]. PooDLe achieves state-of-the-art performance on downstream semantic segmentation and object detection benchmarks, with a notable gain on recognizing small objects. In our ablations, we demonstrate that our joint objective formulation and the SDM are critical for success. In addition, our approach maintains strong performance over a wide range of crop areas and input resolutions while the pooled or dense objective alone quickly degrades
In summary, our contributions are as follows:
  1. 1.
    We introduce PooDLe, a method combining a dense SSL objective based on flow equivariance and a pooled objective using small pseudo-iconic subcrops. PooDLe unifies the two objectives with a spatial decoder module to combine the strengths of both objectives.
  2. 2.
    When pretrained on naturalistic videos, PooDLe achieves state-of-the-art downstream performance on BDD100K [50], ADE20K [53] and Cityscapes [14] semantic segmentation, and BDD100K object detection tasks. PooDLe outperforms prior dense SSL methods [49, 45], particularly on small objects, and is competitive with ImageNet pretraining.
  3. 3.
    We study the effects of crop area, resolution and temporal stride between paired frames. We demonstrate the importance of maintaining pixel density when increasing image resolution for the dense SSL objective. We also perform analysis to show that subcrops can increase prevalence of small objects in the pooled objective by serving as pseudo-iconic views, which supports our experimental findings.
@@
Figure 1: Challenges of SSL on dense, naturalistic video include training on crowded scenes with many objects of varying scale and class imbalance. (a) Iconic image from ImageNet [15]. (b) Dense scene from BDD100K [50]. Global crops (red boxes) typically used in iconic methods may contain disjoint sets of subjects and multiple subjects per crop. Subcrops (green boxes) can provide pseudo-iconic views of a single subject. (c) Long-tailed distribution of classes, by percentage of pixels, from the BDD100K semantic segmentation benchmark. The example subcrops show smaller object classes such as "Traffic Light" and "Person."
@@ 2 Related Work
Representation learning on iconic image datasets has a long history from denoising autoencoders [43] to joint embedding methods [11, 25, 20, 51, 3, 7] to joint-embedding predictive architectures [2, 5]. Joint embedding methods learn representation invariance to visual changes created by augmentations using contrastive [11, 33], mean squared error [20], or classification [7, 6] losses between corresponding pairs, pushing SSL to new heights on ImageNet classification. Later works have extended these methods to internet-scale data [34] and included other modalities like text [37]. Separately, the MAE [26] family of methods learn via reconstruction of masked image regions. iBOT [54] combines joint embedding methods with token reconstruction to produce impressive results on ImageNet classification. [6, 7, 34, 54] all generate small, low-resolution crops (‘multi-crop’) that are optimized to predict the representations of global crops for iconic training examples for little additional compute. This is different from our subcrop strategy which produces aligned image pairs of small crops as pseudo-iconic views from otherwise dense scenes.
@@
Following the success of SSL on ImageNet, other works seek to learn from dense, multi-subject images where augmented views may not contain corresponding subjects for invariance learning.  [45, 48, 8] build upon joint-embedding methods, relying on similarity of features bootstrapped from typical invariance learning to identify positive pairs across dense, unpooled feature maps.  [27, 44] optimize dense losses, contrasting pixels belonging to different semantic classes; these methods require off-the-shelf segmentation modules.  [55, 21] utilize DINO [7] attention maps to identify training pairs, while ADCLR [52] identifies pairs using small ‘query’ crops and the patches that attend to them. All of these methods make strides towards learning from dense images with multiple objects. However, these methods do not consider the effects of a dense objective on per-class accuracy and the skewed class and spatial distributions of more naturalistic data.
@@
Extending beyond images, other works have sought to capture the variance of objects through time by training on pairs of video frames.  Gordon et al. [19] adapts contrastive learning to use correlated frames as positive examples, while [29, 35] identify positive pairs using high representational similarity. FlowE [49] builds on BYOL [20] and identifies positive spatial regions between frames using off-the-shelf flow. Most recently, DoRA [42] proposes a new dense video dataset and extends DINO by clustering over many frames to identify and track objects for representation learning. In the MAE paradigm,  [41, 17] directly reconstruct sequences of frames while [46, 22] perform reconstruction given a corresponding overlapping frame. MC-JEPA [5] learns motion using video data by aligning latent representations throughout the feature pyramid while performing representation learning on ImageNet while VJEPA [4] directly aims to learn video representations via latent video prediction. Like some methods above, PooDLe learns a rich image representation from video data and does so using a dense and pooled objective to capture both scene-level spatial understanding and high-level semantics.
Figure 2: PooDLe combines pooled and dense learning objective for video self-supervised learning. Green path: dense objective performing flow-equivarience learning on the output of the decoder $g(\cdot)$. Orange path: pooled objective encoding $K$ subcrops sampled with flow-informed local cropping. Projector modules are omitted for simplicity. Offline parameter $\xi$ are the exponential moving average (ema) of online parameters $\theta$.
@@ 3 PooDLe[Uncaptioned image]: Pooled and Dense Learning from naturalistic videos
We present PooDLe, a self-supervised method for learning dense visual representations using paired frames from naturalistic, first-person videos. PooDLe combines two SSL objectives: a dense objective for learning representations of dense, crowded scenes; and a pooled objective on small subcrops sampled using flow-aware cropping augmentations. We also propose a lightweight spatial decoder module (SDM) that uses top-down decoder layers and UNet-like lateral connections to earlier encoder representations to both upsample the high-level representations and resurface fine-grained details and small objects that may get lost in downsampling operations. For a high-level overview of PooDLe, see Figure 2.
@@
Inputs to the model are video frame pairs $\bm{x}_{t},\bm{x}_{t+\Delta t}$ with dimensions $H\times W$, and dense optical flow $M_{t\rightarrow t+\Delta t}$ predicted by an external optical flow model. Randomly sampled augmentations $A_{1}$ and $A_{2}$ are applied to each example to create positive training pairs. In a similar setup to BYOL, the encoder and projector are denoted as a function $\bm{p}=f(\bm{x})$ using either online weights $\theta$ or offline, EMA-updated weights $\xi$. The predictor module $q_{\theta}(\cdot)$ only has online weights $\theta$. Separate projector and predictor modules are used for the pooled and dense objectives, but are not annotated for simplicity. We use a ResNet-50 backbone, as well as projectors and predictors following FlowE [49] and BYOL [20], that are discarded after pre-training.
@@
The dense objective follows FlowE [49] by using optical flow $M_{t\rightarrow t+\Delta t}$ to align paired feature projections $\bm{p}_{t}$ and $\bm{p}_{t+\Delta t}$. At a high-level, this objective minimizes differences in representation between corresponding regions. More specifically, the inverse augmentation functions $A^{-1}$ and optical flow are used to align the representations $\bm{p}$ and after upsampling to input resolution $H\times W$, the objective is the squared error:
@@
First we identify $K$ pseudo-iconic subcrop pairs. Unlike for iconic data, random crops from paired frames are unlikely to contain a common subject. To mitigate this, we once again use optical flow in a flow-informed cropping procedure to identify aligned training pairs. For each subcrop pair, we sample a random point $(u,v)$ in the target frame $\bm{x}_{t+\Delta t}$ to serve as the crop center. It is then warped into the earlier frame $\bm{x}_{t}$ using flow $M_{t\rightarrow t+\Delta t}$ plus random jitter $(\delta_{u},\delta_{t})$ for paired center $(u^{\prime},v^{\prime})$. A crop is made around each center, with an area sampled from $U[s_{\text{min}},s_{\text{max}}]$ of the global crop for subcrops $\bm{x}_{t,i}$ and $\bm{x}_{t+\Delta t,i}$.
As we require both crop centers to land within the bounds of the image, subcrops tend to be center-biased [36] and lack diversity. To remedy this, we employ a grid-sampling procedure for selecting the initial crop center $(u,v)$. Each global crop $\bm{x}$ is divided into a grid with cells of side length $d_{\text{grid}}=\text{min}(H,W)\times\sqrt{(s_{\text{min}}+s_{\text{max}})/2}$ for a $H/d_{\text{grid}}\times W/d_{\text{grid}}$ grid. Each cell is selected without replacement and a center $(u,v)$ is then uniformly sampled within the cell.
After $K$ pairs $(x_{t,k},x_{t+\Delta t,k})$ are generated, they are encoded via the backbone and the pooled objective projector. Unlike in the dense objective, no alignment or upsampling is performed, and each projection $\bm{p}$ is averaged-pooled over its spatial dimensions before computing the loss:
@@
where $\bar{\cdot}$ denotes average pooling over spatial dimensions with normalization performed after. Our approach trains each subcrop to predict its corresponding pair, likely containing the same objects in dense scenes. This differs from multi-crop [6], which has local crops predict global crops — potentially less effective for dense scenes where local crops may only capture a subset of objects.
Spatial Decoder Module (SDM).
@@
Figure 3: a) Baseline: both losses combined at final encoder layer; b)PooDLe: SDM incorporates earlier feature maps to upsample for $\mathcal{L}_{\text{dense}}$.
We introduce SDM (Figure 3(b)) to both upsample high-level encoder features and preserve information from lower layers, particularly any smaller foreground objects that may be lost due to pooling operations. Its design draws inspiration from convolutional UNet [38] and FPN [31] and improves upon FlowE’s use of dilated convolutions to replace pooling by efficiently maintaining high-resolution representations while significantly reducing activations and memory usage.
@@
$\bm{z}^{l+1}=g(upsample(\bm{z}^{(l)}))+lateral(\bm{z}^{j}),$ (3)
where $\bm{z}^{(l)}$ is the representation after the $l^{\text{th}}$ encoder stage and $\bm{z}^{(j)}$ is an earlier feature map of the same spatial dimensions as $\bm{z}^{(l+1)}$. The use of computation blocks and lateral connections are ablated in Table 4.2. Figure 3 depicts a naive implementation that places both objectives at the top encoder level and PooDLe, which uses the SDM to integrate the two objectives in a complementary fashion.

### v2

@@ Abstract
Self-supervised learning has driven significant progress in learning from single-subject, iconic images. However, there are still unanswered questions about the use of minimally-curated, naturalistic video data, which contain dense scenes with many independent objects, imbalanced class distributions, and varying object sizes. In this paper, we propose PooDLe, a self-supervised learning method that combines an invariance-based objective on pooled representations with a dense SSL objective that enforces equivariance to optical flow warping. Our results show that a unified objective applied at multiple feature scales is essential for learning effective image representations from naturalistic videos. We validate our method with experiments on the BDD100K driving video dataset and the Walking Tours first-person video dataset, demonstrating its ability to capture spatial understanding from a dense objective and semantic understanding via a pooled representation objective.
@@ 1 Introduction
Humans and other animals learn visual understanding from a continuous stream of inputs with little explicit supervision. Self-supervised learning (SSL) (Chen et al., 2020; Grill et al., 2020; Chen & He, 2021; Caron et al., 2021; Bardes et al., 2021; He et al., 2022; Assran et al., 2023; Bardes et al., 2023a; He et al., 2020) has made great strides in learning without human annotations, becoming competitive with supervised learning. However, many methods still revolve around ImageNet (Deng et al., 2009), which is implicitly supervised through iconic images that contain a single, clear subject and a balanced class distribution. In contrast, naturalistic data like egocentric videos contain cluttered scenes, imbalanced classes, and objects of varying sizes, making them ill-suited for iconic methods.
@@
Nevertheless, these naturalistic videos are still valuable for their information density and ease of collection, while also mimicking the real-life perspective of humans. Unfortunately, iconic methods, which pool global image representations, may perform poorly as dense scenes produce views containing independent subjects that are semantically incompatible (Figure 1, red boxes). Recent works have attempted to address this weakness by introducing 1) cropping (Selvaraju et al., 2021) or attention (Venkataramanan et al., 2024) mechanisms to account for multiple subjects, and 2) “dense SSL” objectives (Xiong et al., 2021; Wang et al., 2021b) with losses defined over regions of unpooled image representations.
@@
While dense SSL methods avoid semantic mismatches, we discover that they are susceptible to spatial imbalance where larger background classes like the sky dominate the representation, while smaller classes like pedestrians are underrepresented. This is undesirable because smaller foreground objects should be prioritized over low-detail, repetitive background classes. Furthermore, this can be dangerous in applications like self-driving (Yu et al., 2020) where critical objects like pedestrians occupy less than $0.3\%$ of a video frame (Figure 1, green boxes and 1). This contrasts with ImageNet (Deng et al., 2009) training, where models can easily learn semantics from iconic images with clear, single-subject views and a balanced class distribution. Surprisingly, dense methods like FlowE (Xiong et al., 2021) and supervised ImageNet pretraining achieve similar downstream performance while converging to very different solutions; the former prioritizes large background classes while the latter captures many small and rare classes, but with relatively poor specificity. Some dense SSL methods (Wang et al., 2021b; Parthasarathy et al., 2023; Venkataramanan et al., 2024) include losses that optimize a global, pooled representation to learn semantic information from dense scenes, but do not explore how to integrate these two objectives through means of architecture and augmentation strategies.
@@
Figure 1: Challenges of SSL on dense, naturalistic video include training on crowded scenes with many objects of varying scale and class imbalance. (a) Iconic image from ImageNet. (b) Dense scene from BDD100K. Global crops (red boxes) used for iconic data contain multiple, disjoint sets of objects. Subcrops (green boxes) provide pseudo-iconic views of a single subject. (c) Naturalistic video like BDD100K have a long-tailed class distribution of class sizes and frequency; classes ordered by percentage of pixels. Example subcrops capture smaller classes like "Traffic Light" and "Person."
To address the challenges of cluttered scenes and spatial imbalance, we propose a joint Pooled and Dense Learning (PooDLe) method that optimizes a dense SSL objective over full images and a pooled objective on smaller, semantically-aligned views. The combination of objectives captures both high-level semantics and fine-grained details, effectively representing both small objects and scene-level understanding. Our dense objective is adopted from FlowE, using optical flow warping to align dense feature maps. To adapt the pooled objective to dense scene data, we introduce a flow-informed cropping procedure that generates pairs of smaller “subcrops” with high alignment. These subcrops serve as pseudo-iconic views of foreground objects, functionally increasing the prevalence of smaller objects (Figure 1). Finally, we introduce a lightweight spatial decoder module (SDM) with top-down layers and UNet-like lateral connections (Ronneberger et al., 2015) to upsample high-level semantic representations and preserve smaller objects in the dense objective. We show that both objectives, combined with the SDM, is essential for capturing the semantics of smaller objects and achieving strong downstream task performance.
We pretrain on BDD100K, a dataset of dashcam driving videos as well as Walking Tours, a dataset of first-person walking videos (Venkataramanan et al., 2024). PooDLe achieves state-of-the-art performance on semantic segmentation and object detection benchmarks, with a notable gain on recognizing small objects. We also introduce Walking Tours Semantic (WT-Sem) as a new in-distribution semantic segmentation evaluation for Walking Tours. In our ablations, we show that our joint objective formulation and the SDM are critical for success. Finally, we study the effect of crop area, input resolution, number of subcrops and temporal stride.
In summary, our contributions are as follows:
  1. 1.
    We introduce PooDLe, a new SSL method which overcomes the challenges of spatial imbalance and cluttered scenes by unifying a flow equivariance, dense SSL objective and a pooled objective over pseudo-iconic subcrops alongside a spatial decoder module to effectively learn from naturalistic video.
  2. 2.
    We deconstruct the BDD100K semantic segmentation task, identifying class categories by frequency and size within the dataset. We show that existing dense SSL methods and supervised ImageNet training produce different results across these categories, while PooDLe learns a balanced semantic and spatial representation to achieve strong, consistent performance.
  3. 3.
    We study the effects of global and subcrop area, input resolution and temporal stride between paired frames.
@@ 2 Related Work
Representation learning on iconic image datasets has a long history, from denoising autoencoders (Vincent et al., 2010) to joint embedding methods (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Zbontar et al., 2021; Bardes et al., 2021; Caron et al., 2021) to joint-embedding predictive architectures (Assran et al., 2023; Bardes et al., 2023b). Joint embedding methods learn representation invariance to visual changes produced by augmentations using contrastive (Chen et al., 2020; Oord et al., 2018), mean squared error (Grill et al., 2020), or classification (Caron et al., 2021; 2020) losses between corresponding pairs, pushing SSL to new heights on ImageNet classification. Later works extend these methods to curated, internet-scale data (Oquab et al., 2023) and include other modalities like text (Radford et al., 2021). Separately, MAE (He et al., 2022) learns via reconstruction of masked image regions. iBOT (Zhou et al., 2021) combines joint embedding methods with token reconstruction to achieve impressive results on ImageNet classification. The methods above have been primarily designed for iconic images and contain assumptions that may not transfer well to uncurated datasets, e.g. dense scenes. Methods leveraging multi-crop (Caron et al., 2020; 2021; Oquab et al., 2023; Zhou et al., 2021) generate small crops optimized to predict the representations of global crops for training on iconic images with little additional compute. In contrast, our subcrop strategy yields small, aligned crops as pseudo-iconic, paired views from otherwise dense scenes.
@@
Following the success of SSL on ImageNet, other works seek to learn from dense, multi-subject images where augmented views may not contain corresponding subjects for invariance learning. Wang et al. (2021b); Xie et al. (2021); Chen et al. (2021) extend joint embedding methods by leveraging feature similarity bootstrapped from standard invariance learning to identify positive pairs across dense, unpooled feature maps. Hénaff et al. (2021); Wang et al. (2021a) optimize dense losses, contrasting pixels belonging to different semantic classes; these methods require off-the-shelf segmentation modules. Ziegler & Asano (2022); Guo et al. (2023) utilize DINO (Caron et al., 2021) attention maps to identify training pairs, while ADCLR (Zhang et al., 2023b) identifies pairs using small “query” crops and the patches that attend to them. These methods advance the ability to learn from dense images with multiple objects, but still have limitations. Some rely on learning objectives that make assumptions about iconic data, while others struggle with the spatial imbalance problem that is especially prevalent in naturalistic data.
@@
Extending beyond images, other works have sought to capture the variance of objects through time by training on pairs of video frames. Gordon et al. (2020) adapts contrastive learning to use correlated frames as positive examples, while Jabri et al. (2020); Parthasarathy et al. (2023) identify positive pairs based on high similarity in representation space. FlowE (Xiong et al., 2021) builds on BYOL (Grill et al., 2020) and identifies positive spatial regions between frames using off-the-shelf flow. MC-JEPA (Bardes et al., 2023b) learns motion using video data by aligning latent representations throughout the feature pyramid while performing representation learning on ImageNet. Most recently, DoRA (Venkataramanan et al., 2024) proposes a new dense video dataset and extends DINO by clustering over many frames to identify and track objects for representation learning. In the MAE paradigm, Tong et al. (2022); Feichtenhofer et al. (2022) directly reconstruct sequences of frames while Weinzaepfel et al. (2022); Gupta et al. (2023) perform reconstruction given a corresponding overlapping frame. While PooDLe learns a rich image representation from video data similar to these existing methods, it distinguishes itself by leveraging a unified dense and pooled objective architecture, specifically designed to tackle the challenges posed by naturalistic data.
Figure 2: PooDLe, a self-supervised learning method that combines pooled and dense objectives. Green path: dense objective performing flow-equivariance learning on the output of the decoder $g(\cdot)$. Orange path: pooled objective encoding $K$ subcrops sampled with flow-informed cropping. Projector modules are not shown. Offline weights $\xi$ are the exponential moving average of online weights $\theta$.
@@ 3 PooDLe[Uncaptioned image]: Pooled and Dense Learning from naturalistic videos
We present PooDLe, a self-supervised method for learning visual representations using paired frames from naturalistic, first-person videos. PooDLe combines two SSL objectives: a dense objective for learning representations of dense, crowded scenes; and a pooled objective on small subcrops sampled using flow-aware cropping augmentations. We also propose a lightweight spatial decoder module (SDM) that uses top-down decoder layers and UNet-like lateral connections to earlier encoder representations to both upsample the high-level representations and resurface fine-grained details and small objects that may get lost in downsampling operations. For a high-level overview of PooDLe, see Figure 2.
@@
Inputs to the model are video frame pairs $\bm{x}_{t},\bm{x}_{t+\Delta t}$ with dimensions $H\times W$, and dense optical flow $M_{t\rightarrow t+\Delta t}$. Randomly sampled augmentations $A_{1}$ and $A_{2}$ are applied to each example to create positive training pairs. In a similar setup to BYOL, the encoder and projector are denoted as a function $\bm{p}=f(\bm{x})$ using either online weights $\theta$ or offline, EMA-updated weights $\xi$. The predictor module $q_{\theta}(\cdot)$ only has online weights $\theta$. Separate projector and predictor modules are used for the pooled and dense objectives, but are not annotated for simplicity. We use a ResNet-50 backbone, as well as projectors and predictors following FlowE (Xiong et al., 2021) and BYOL (Grill et al., 2020), that are discarded after pretraining.
@@
The dense objective follows FlowE (Xiong et al., 2021) by using optical flow $M_{t\rightarrow t+\Delta t}$ to align paired feature projections $\bm{p}_{t}$ and $\bm{p}_{t+\Delta t}$. At a high-level, this objective minimizes differences in representation between corresponding regions. More specifically, the inverse augmentation functions $A^{-1}$ and optical flow are used to align the representations $\bm{p}$ and after upsampling to input resolution $H\times W$, the objective is the squared error:
@@
First, we identify $K$ pseudo-iconic subcrop pairs. Unlike for iconic data, random crops from paired frames are unlikely to contain a common subject. To mitigate this problem, we once again use optical flow in a flow-informed cropping procedure to identify aligned training pairs. For each subcrop pair, we sample a random point $(u,v)$ in the target frame $\bm{x}_{t+\Delta t}$ to serve as the crop center. It is then warped into the earlier frame $\bm{x}_{t}$ using flow $M_{t\rightarrow t+\Delta t}$ plus random jitter $(\delta_{u},\delta_{t})$ for paired center $(u^{\prime},v^{\prime})$. A crop is made around each center, with an area sampled from $U[s_{\text{min}},s_{\text{max}}]$ of the global crop for subcrops $\bm{x}_{t,i}$ and $\bm{x}_{t+\Delta t,i}$.
As we require both crop centers to land within the bounds of the image, subcrops tend to be center-biased (Peng et al., 2022) and lack diversity. To remedy this, we employ a grid-sampling procedure for selecting the initial crop center $(u,v)$. Each global crop $\bm{x}$ is divided into a grid with cells of side length $d_{\text{grid}}=\text{min}(H,W)\times\sqrt{(s_{\text{min}}+s_{\text{max}})/2}$ for a $H/d_{\text{grid}}\times W/d_{\text{grid}}$ grid. Each cell is selected without replacement, and a center $(u,v)$ is then uniformly sampled within the cell.
After $K$ pairs $(x_{t,k},x_{t+\Delta t,k})$ are generated, they are encoded by the backbone and the pooled objective projector. Unlike the dense objective, no alignment or upsampling is performed, and each projection $\bm{p}$ is averaged-pooled over its spatial dimensions before computing the loss:
@@
where $\bar{\cdot}$ denotes average pooling over spatial dimensions followed by normalization. Our objective has each subcrop to predict its corresponding pair, which contains the same object in a different frame. This differs from multi-crop (Caron et al., 2020), where local crops predict global crops, which would be less effective for dense scenes as local crops only capture a subset of the objects in a frame.
@@
Figure 3: a) Baseline: both losses combined at final encoder layer; b)PooDLe: SDM incorporates earlier feature maps, upsampling for $\mathcal{L}_{\text{dense}}$.
Spatial Decoder Module (SDM).
We introduce the SDM (Figure 3(b)) to upsample high-level encoder features and preserve information from lower layers, particularly smaller foreground objects that may be lost during pooling operations. Its design draws inspiration from a convolutional UNet (Ronneberger et al., 2015) and FPN (Lin et al., 2017) and improves upon FlowE’s use of dilated convolutions to replace pooling by efficiently maintaining high-resolution representations while reducing activations and memory usage.
@@
$\bm{z}^{l+1}=g(\text{upsample}(\bm{z}^{(l)}))+\text{lateral}(\bm{z}^{j}),$ (3)
where $\bm{z}^{(l)}$ is the representation after the $l^{\text{th}}$ encoder stage and $\bm{z}^{(j)}$ is an earlier feature map of the same spatial dimensions as $\bm{z}^{(l+1)}$. The use of computation blocks and lateral connections is ablated in Table 4.2. Figure 3 contrasts a naive implementation that places both objectives at the top encoder level and PooDLe, which uses the SDM to integrate the two objectives in a complementary fashion.